OpenStudy (anonymous):

A 10-foot ladder leans against a wall with its foot braced 3 feet from wall’s base. How far up the wall does the ladder reach?

jimthompson5910 (jim_thompson5910):

Hint: 

jimthompson5910 (jim_thompson5910):

use the pythagorean theorem to solve for x a^2 + b^2 = c^2 x^2 + 3^2 = 10^2 x^2 + 9 = 100 I'll let you finish

OpenStudy (anonymous):

\[100-9= 81 =\sqrt{9} ?\]

OpenStudy (anonymous):

9.53

OpenStudy (anonymous):

>.< Dang it, how?

OpenStudy (anonymous):

I did reverse Pythagorean theorem \[c ^{2} - b ^{2} = a ^{2}\]

jimthompson5910 (jim_thompson5910):

\[\Large x^2 + 9 = 100\] \[\Large x^2 = 100-9\] \[\Large x^2 = 91\] \[\Large x = \sqrt{91}\] \[\Large x \approx 9.53939201416946\]

OpenStudy (anonymous):

height = sqrt(100 - 9 ) = sqrt (91) approx 9.53

OpenStudy (anonymous):

I was told to leave it in radical form

OpenStudy (anonymous):

oh well yeah 9 then

OpenStudy (anonymous):

Yay! Thank you everyone

jimthompson5910 (jim_thompson5910):

so leave it as \[\Large \sqrt{91}\]

OpenStudy (anonymous):

Just to let it be known I realized that I was wrong. I just realized that I got 81 and it should have been 91, then I finished the problem and it came out to be what you guys got , 9.53 .. thank you for the help [: